OpenStudy (anonymous):

Let Y be the number of requests, per minute for files from a file server. Y is a random variable with Poisson distribution with parameter 6. Let \(Y_1, Y_2, Y_3\) denote the number of requests in 3 successive minutes. Assume that \(Y_i\) have same distribution. Let for each \(i = 1, 2, 3, A_i\) be the event \(\{Y_i=0\}\). Assume also that the events \(A_1, A_2, A_3\) are mutually independent. a) Calculate the probability that there will be no requests in entire 3 minutes period. b) Give a full set of conditions that ensure that \(A_1,A_2,A_3\) are mutually independent. chars limit ended

OpenStudy (anonymous):

a) So answer should be just \(P(Y_1=0)+P(Y_2=0)+P(Y_3=0)=3\cdot e^{-6}\) ? b) What does mutually independent means? is it mutual exclusive event or independent events?

OpenStudy (anonymous):

What is a Poisson distribution?

OpenStudy (anonymous):

same as binomial distribution but n approaches to infinity

OpenStudy (jamesj):

For a, you use the same procedure as we do for all such indep probabilities: multiply the three probabilities, don't add.

OpenStudy (jamesj):

A set of events { Pj } are mutually independent if \[ P(P_j | P_k) = P_j \ \ \hbox{ for all } j \neq k \]

OpenStudy (anonymous):

sorry. i am trying to figure out what this is really asking. \[Y_1=0\] means no calls in the first minute?

OpenStudy (anonymous):

do I need to prove it or just enough to write \(P(Y1|Y2)=P(Y1)\) (and others) since I am asked to give full set of conditions?

OpenStudy (anonymous):

i think yes

OpenStudy (anonymous):

and therefore \[P(Y_1=0)=e^{-3}\] yes?

OpenStudy (jamesj):

@tomas.a, you can't "prove it"; you're basically writing down a careful definition of what mutually indep here means.

OpenStudy (anonymous):

\(e^{-6}\)

OpenStudy (anonymous):

sorry typo there, right \[P(Y_1=0)=e^{-6}\]

OpenStudy (anonymous):

pmf for Poisson is \(P(x)=\frac{\lambda^x e^{-\lambda}}{x!}\)

OpenStudy (anonymous):

@jamesj correct me if i am wrong, but if this is poisson with mean 6 this means expect 6 calls per minute, and also necessarily 18 calls every 3 minutes. meaning that the probability you get no calls in 3 successive minutes must be \[e^{-18}\]

OpenStudy (anonymous):

yeah he said that I need to multiply, not add so \(e^{-6}\cdot e^{-6}\cdot e^{-6}=e^{-18}\)

OpenStudy (anonymous):

which by the laws of exponents is exactly what you get if you multiply the probabilities together

OpenStudy (jamesj):

Exactly.

OpenStudy (anonymous):

yes you multiply because 1) they are independent and 2) it is an AND

OpenStudy (anonymous):

statement

OpenStudy (anonymous):

but i am confused by part 2 because if i remember correctly the fact that you have a poisson distribution means they MUST be independent

OpenStudy (anonymous):

yeah i know that AND multiplication and OR addition but i mixed them ><